# Introducing NewCluster: the first half of the history of a high-resolution cluster simulation

### v1

@@ Abstract
Methods. The zoom-in simulation targets a volume of $4.1\sigma$ overdensity region, which is expected to evolve into a galaxy cluster with a virial mass of $5\times 10^{14}M_{\odot}$, comparable to that of the Virgo Cluster. The zoom-in volume extends out to 3.5 virial radii from the central halo. The novelties of NewCluster are found in its resolutions. Its stellar mass resolution of $2\times 10^{4}M_{\odot}$ is effective for tracing the early assembly of massive galaxies as well as the formation of dwarf galaxies. The spatial resolution of 68 parsec in the best-resolved regions in the adaptive-mesh-refinement approach is powerful to study the detailed kinematic structure of galaxies. The time interval between snapshots is also exceptionally short—15 Myr—which is ideal for monitoring changes in the physical properties of galaxies, particularly during their orbital motion within a larger halo. The simulation has up-to-date feedback schemes for supernovae and active galactic nuclei. The chemical evolution is calculated for ten elements, along with dust calculation that includes the formation, size change, and destruction. To overcome the limitations of the Eulerian approach used for gas dynamics in this study, we employ Monte Carlo-based tracer particles in NewCluster, enabling a wide range of scientific investigations.
@@ 1 Introduction
Numerical simulations have become an indispensable tool in modern astronomy for exploring the complex and non-linear nature of the Universe. A pioneering study by Toomre & Toomre (1972), which modeled the morphologies of interacting galaxies, demonstrated the potential of numerical simulations to capture such non-linearity. Since then, continuous advancements in computational methods have established numerical methods as an important complement to observational and theoretical approaches (e.g., White & Rees, 1978; Davis et al., 1985; Hernquist, 1987; Cole, 1991; White & Frenk, 1991; Kauffmann et al., 1993).
@@
These early successes have laid the foundation for the advancement of contemporary cosmological simulations. The cosmological principle assumes that both the spatial distribution of matter and its power spectrum approach homogeneity on sufficiently large scales. This justifies the use of a large numerical volume to approximate a representative subregion of the Universe. Provided this, cosmological simulations are initialised with primordial density fluctuations consistent with the early Universe and evolved forward in time using the governing equations of gravity and hydrodynamics, coupled with astrophysical subgrid models. These calculations yield the formation and evolution of self-consistent large-scale structure within the simulated volume (e.g., Peebles, 1982; Davis et al., 1985; White et al., 1987).
Observational validation of the $\Lambda$CDM cosmology, including precise measurements of cosmological parameters (e.g., Komatsu et al., 2011), significantly accelerated the development of early cosmological numerical simulations. A prominent example is the Millennium simulation (Springel et al., 2005), based on the $\Lambda$CDM model, which successfully reproduced large-scale structures consistent with that observed in the survey (Springel et al., 2006). This achievement sets a stage for practically attainable theoretical investigation of nonlinear cosmological structures. Although the simulation itself included only collisionless dark matter (DM) particles, subsequent studies incorporated semi-analytic models to populate galaxies, demonstrating an important link between the non-linear nature of the Universe and galaxy evolution (e.g., Bower et al., 2006; De Lucia & Blaizot, 2007).
@@
The next generation of cosmological simulations incorporated hydrodynamics to model baryonic physics, including cooling and heating, star formation, and feedback, through dedicated astrophysical prescriptions. Seminal large-volume simulations that pioneered this approach include Eagle (Schaye et al., 2015), Horizon-AGN (Dubois et al., 2014a), and Illustris (Vogelsberger et al., 2014b). These simulations typically achieved spatial resolutions of $\sim 1\,{\rm kpc}$, sufficient to capture the global properties and statistical trends of galaxy populations in cosmological contexts (Vogelsberger et al., 2014a; Furlong et al., 2015; Dubois et al., 2016). The larger volumes further allowed for systematic studies of environmental effects on galaxy evolution by treating environment as a controllable parameter (e.g., Jeon et al., 2022). However, a resolution of $\sim 1\,{\rm kpc}$ remains insufficient to resolve the internal structure and processes with sub-kpc physical scales within galaxies. For example, in a Milky Way-like galaxy with a size of $\sim 10\,{\rm kpc}$, only a few resolution elements are available to represent its internal structure.
@@
Recent cosmological simulations have reached the sub-kpc resolution (e.g., Tremmel et al., 2017; Pillepich et al., 2019). With this improvement in resolution, galactic substructures begin to be resolved in a cosmological context (e.g., Du et al., 2021), enabling detailed analyses of the kinematic and spatial properties of galactic components. These sub-kpc simulations have also resolved structures beyond galaxies themselves, such as circumgalactic medium (CGM) (e.g., Peeples et al., 2019), making it feasible to study the baryon cycle of galaxies. However, despite these advances, it is sometimes insufficient to resolve finer galactic structures, such as the multiphase nature of interstellar medium (ISM) or the thin disc of spiral galaxies.
@@
Subsequent cosmological hydrodynamic simulations have pushed the resolution below $\sim 100\,{\rm pc}$ (e.g., Wang et al., 2015; Wetzel et al., 2016; Dubois et al., 2021; Yi et al., 2024), with some achieving resolutions as fine as sub-10 pc (Hopkins et al., 2018b). This level of spatial resolution enables resolving both the distinction of thin and thick discs of galaxies (e.g., Park et al., 2021; Renaud et al., 2021; Yi et al., 2024) or compact systems (Jang et al., 2024) and the multiphase structure of galactic gas, including the warm and cold ISM (Dubois et al., 2021). Such high resolution has facilitated the implementation of more advanced subgrid models. Mechanical supernova feedback, which has been reported to be more effective in suppressing star formation in low-mass galaxies, has been implemented in several high-resolution simulations, enabled by the ability to assign momentum to gas elements in the immediate vicinity of star particles (Kimm & Cen, 2014; Hopkins et al., 2018a). Direct measurements of the differential velocity within the ISM have been used to implement the gravo-thermo-turbulent model of star formation efficiency (Kimm et al., 2017; Kretschmer & Teyssier, 2020), allowing detailed investigations on star formation activity within galaxies (Kraljic et al., 2024).
@@
Furthermore, the corresponding increase in baryonic (and DM) mass resolution is also important, as poor resolutions introduce numerical biases (e.g., Klypin et al., 1999; Ludlow et al., 2019). Fine mass resolutions allow for the detection of structures with extremely low surface brightness, such as intracluster light (ICL), ultra-diffuse galaxies (UDGs), tidal streams, and ultra-faint dwarfs. A stellar particle mass of $\sim 10^{4}\,\,M_{\sun}$ enables the simulation to reach surface brightness levels as faint as $34\,{\rm mag\,arcsec^{-2}}$ at $z=0.1$. This is enough to reproduce results from deep imaging surveys such as Large Synoptic Survey Telescope (LSST), Euclid, and James Webb Space Telescope (JWST) (Euclid Collaboration et al., 2022).
@@
Therefore, more realistic and advanced cosmological simulations at the current stage require a sub-100 pc resolution and a high mass resolution. A major challenge arising from these requests is the high computational cost due to the enormous number of elements and associated calculations. A good bypass to satisfy the needs above and to avoid this difficulty is running a simulation for a small interested region with high resolutions, while the rest of the volume is treated with relatively poorer resolutions: that is, a zoom-in approach (e.g., Navarro & White, 1994; Barnes et al., 2017; Choi & Yi, 2017; Hopkins et al., 2018b; Tremmel et al., 2019; Dubois et al., 2021).
@@
The majority of recent high-resolution (sub-100 pc) zoom-in simulations focus on individual galaxies or average-density fields of the Universe. This has resulted in a lack of representative high-resolution counterparts for dense environments such as galaxy clusters. Galaxy clusters are thought to have formed from the growth of primordial over-density in the early Universe (Kravtsov & Borgani, 2012), implying that developing clusters were once among the most active sites of star formation (Jeon et al., 2022). Therefore, proto-clusters are the primary candidates for the sites for containing high-redshift galaxies that are frequently detected by the JWST (Labbé et al., 2023; Carniani et al., 2024).
@@
In this regard, we carry out the NewCluster simulation, a high-resolution cosmological hydrodynamic simulation designed to reproduce the formation and evolution of a galaxy cluster down to near-local redshift. NewCluster serves as a massive halo counterpart to existing high-resolution zoom-in simulations on average-density fields, NewHorizon and NewHorizon2, particularly because it shares many of its subgrid prescriptions with them. The simulation aims to monitor the formation and evolution of a massive cluster and galaxy evolution inside the extreme environment. Moving at $\sim 10^{3}\,{\rm km\,s^{-1}}$, galaxies experience rapid gas stripping through hydrodynamic ram pressure and DM and stellar stripping through gravitational tidal interactions (Smith et al., 2016; Rhee et al., 2017; Han et al., 2018; Jung et al., 2018). Such processes are likely responsible for jellyfish (Chung et al., 2009) and perhaps ultra-diffuse galaxies (van Dokkum et al., 2015). We also expect the transformation of the morphology of galaxies inside the cluster halo, while the impact of the cluster merger (starting at $z=0.8$) is open to wild speculations.
@@
This paper is organised as follows. In Sect. 2, we describe the simulation code, astrophysical prescriptions, and post-processing methods. In Sect. 3, we present a general overview of the simulation and the galaxy scaling relations at $z=0.8$, along with key examples of applications of the cluster simulation to various topics. In Sect. 4, we summarise the main findings of the paper.
@@ 2.1 Code: RAMSES-yOMP
RAMSES (Teyssier, 2002) is an astrophysical simulation code based on the adaptive mesh refinement (AMR) technique. In this study, we use a specialised branch of the code, called RAMSES-yOMP (Han et al., 2025), which has been extensively modified to enhance simulation performance. The improvement is primarily due to the implementation of a hybrid parallelization scheme that utilises a dual-level multiprocessing structure to enhance the parallel scalability, combining the message passing interface (MPI) and open multi-processing (OMP) libraries. A series of modifications to yOMP are described in Han et al. (2025), including updates on subroutines such as the load-balancing scheme and the conjugate gradient Poisson solver, aimed at improving the overall performance of the simulation. Load balancing is enhanced by assigning greater computational weights to particles that require more processing, particularly those involved in complex feedback and density estimation. In a test suite simulating a cosmological volume with 1,536 computing cores, the modified code achieved approximately a two times speed-up compared to the pre-modified version.
@@ 2.2 Initial Condition and Zoom-in Region
The initial condition is generated using MUSIC (Hahn & Abel, 2011) at $z=50$ with the WMAP7 cosmology (Komatsu et al., 2011) with parameters $\Omega_{\Lambda}=0.728$, $\Omega_{m}=0.272$, $\sigma_{8}=0.810$, $n_{s}=0.967$, $h_{0}=0.703$ and $\Omega_{b}=0.0455$. The transfer function from Eisenstein & Hu (1998) and CAMB (Lewis et al., 2000) is used to derive the initial distribution of DM and baryons, respectively. The zoom-in technique is utilised to efficiently simulate cluster regions in high resolution. First, we ran a DM-only simulation in a box with a volume of $(100\,h^{-1}\,\mathrm{Mpc})^{3}$ with periodic boundary conditions. We use the Rockstar halo finder (Behroozi et al., 2013b) to detect halos within the volume at $z=0$ and select a cluster with a virial mass of $4.7\times 10^{14}\,M_{\sun}$. A spherical region centred on the density peak of the target cluster is set with a radius of $3.5R_{\rm vir}$, where $R_{\rm vir}$ refers to the virial radius [^1] In Rockstar, the virial radius is defined the range within which contains $\Delta_{\rm vir}(z)\rho_{c}$ (Bryan & Norman, 1998; Behroozi et al., 2013b). of the target cluster, to record the identity of the confined particles and trace their initial distribution at $z=50$. The zoom-in region, corresponding to a $4.1\sigma$ overdensity above the Universal mean, is defined as a convex hull generated based on the distribution of the recorded particles in the initial volume. The selected convex hull roughly covers a comoving volume of $23,300\,\mathrm{Mpc}^{3}$ with a total mass of $8.6\times 10^{14}\,M_{\sun}$.
@@
The sudden increase in resolution at the specific scale factors may introduce numerical artifacts in the gas properties of galaxies. To mitigate this, we smooth out the transitions at which refinement is applied. We follow the approach of Snaith et al. (2018), but additionally apply the same method to the Jeans length criterion. During the transition phase, we continuously varied the parameter $N_{\rm jeans}$ from 0 to 1 and $m_{\rm refine}$ from 100 to 8, at the transition phase to perform smooth refinement, i.e.,
@@
where $a$ is the current scale factor, $a_{0}$ is the centre of the transition phase (defined as 0.1, 0.2, 0.4, 0.8 at grid levels 18, 19, 20, 21), and $w$ is the half-width of the transition phase, set to $0.025$ in this work. $C$ is the refinement threshold and can be either $N_{\rm Jeans}$ or $m_{\rm refine}$, where $C_{L}$ and $C_{R}$ denote the initial and final values for the refinement transition, respectively. The parameter pair $[C_{L},C_{R}]$ is set to $[0,1]$ for $N_{\rm Jeans}$, and $[100,8]$ for $m_{\rm refine}$. $m_{\rm refine}$ defines the mass increase required for a cell to be refined. For example, if $m_{\rm refine}=8$, the mass within a cell needs to be eight times higher than the initial cell mass to be eligible for refinement. $N_{\rm Jeans}$ is the parameter for the Jeans length refinement criterion, in which the cell with a hydrogen gas density $n\geq 5\,\rm H\,cm^{-3}$ can be refined if

### v2

@@ Abstract
Methods. The zoom-in simulation targets a volume of $4.1\sigma$ overdensity region, which is expected to evolve into a galaxy cluster with a virial mass of $5\times 10^{14}M_{\odot}$, comparable to that of the Virgo Cluster. The zoom-in volume extends out to 3.5 virial radii from the central halo. The novelties of NewCluster are found in its resolutions. Its stellar mass resolution of $2\times 10^{4}M_{\odot}$ is effective for tracing the early assembly of massive galaxies as well as the formation of dwarf galaxies. The spatial resolution of 68 parsecs in the best-resolved regions in the adaptive-mesh-refinement approach is powerful to study the detailed kinematic structure of galaxies. The time interval between snapshots is also exceptionally short—15 Myr—which is ideal for monitoring changes in the physical properties of galaxies, particularly during their orbital motion within a larger halo. The simulation has up-to-date feedback schemes for supernovae and active galactic nuclei. The chemical evolution is calculated for ten elements, along with dust calculation that includes the formation, size change, and destruction. To overcome the limitations of the Eulerian approach used for gas dynamics in this study, we employ Monte Carlo-based tracer particles in NewCluster, enabling a wide range of scientific investigations.
@@ 1 Introduction
Numerical simulations have become an indispensable tool in modern astronomy for exploring the complex and non-linear nature of the Universe. A pioneering study by TT72, which modeled the morphologies of interacting galaxies, demonstrated the potential of numerical simulations to capture such non-linearity. Since then, continuous advancements in computational methods have established numerical methods as an important complement to observational and theoretical approaches (e.g., WR78; Davis85; Hernquist87; Cole91; WF91; Kauffmann93).
@@
These early successes have laid the foundation for the advancement of contemporary cosmological simulations. The cosmological principle assumes that both the spatial distribution of matter and its power spectrum approach homogeneity on sufficiently large scales. This justifies the use of a large numerical volume to approximate a representative subregion of the Universe. Provided this, cosmological simulations are initialised with primordial density fluctuations consistent with the early Universe and evolved forward in time using the governing equations of gravity and hydrodynamics, coupled with astrophysical subgrid models. These calculations yield the formation and evolution of self-consistent large-scale structure within the simulated volume (e.g., Peebles82; Davis85; White87).
[…]
@@
The next generation of cosmological simulations incorporated hydrodynamics to model baryonic physics, including cooling and heating, star formation, and feedback, through dedicated astrophysical prescriptions. Seminal large-volume simulations that pioneered this approach include Eagle (Schaye15), Horizon-AGN (2014MNRAS.444.1453D), and Illustris (Vogelsberger14). These simulations typically achieved spatial resolutions of $\sim 1\,{\rm kpc}$, sufficient to capture the global properties and statistical trends of galaxy populations in cosmological contexts (Vogelsberger14b; Furlong15; Dubois16). The larger volumes further allowed for systematic studies of environmental effects on galaxy evolution by treating environment as a controllable parameter (e.g., 2022ApJ...941....5J). However, a resolution of $\sim 1\,{\rm kpc}$ remains insufficient to resolve the internal structure and processes with sub-kpc physical scales within galaxies. For example, in a Milky Way-like galaxy with a size of $\sim 10\,{\rm kpc}$, only a few resolution elements are available to represent its internal structure.
@@
Recent cosmological simulations have reached the sub-kpc resolution (e.g., Tremmel17; Pillepich19). With this improvement in resolution, galactic substructures begin to be resolved in a cosmological context (e.g., Du21), enabling detailed analyses of the kinematic and spatial properties of galactic components. These sub-kpc simulations have also resolved structures beyond galaxies themselves, such as circumgalactic medium (CGM) (e.g., Peeples19), making it feasible to study the baryon cycle of galaxies. However, despite these advances, it is sometimes insufficient to resolve finer galactic structures, such as the multiphase nature of interstellar medium (ISM) or the thin disc of spiral galaxies.
@@
Subsequent cosmological hydrodynamic simulations have pushed the resolution below $\sim 100\,{\rm pc}$ (e.g., Wang15; Wetzel16; 2021A&A...651A.109D; Yi24), with some achieving resolutions as fine as sub-10 pc (Hopkins18b). This level of spatial resolution enables resolving both the distinction of thin and thick discs of galaxies (e.g., Park21; renaud2021; Yi24) or compact systems (Jang2024) and the multiphase structure of galactic gas, including the warm and cold ISM (2021A&A...651A.109D). Such high resolution has facilitated the implementation of more advanced subgrid models. Mechanical supernova feedback, which has been reported to be more effective in suppressing star formation in low-mass galaxies, has been implemented in several high-resolution simulations, enabled by the ability to assign momentum to gas elements in the immediate vicinity of star particles (2014ApJ...788..121K; Hopkins18a). Direct measurements of the differential velocity within the ISM have been used to implement the gravo-thermo-turbulent model of star formation efficiency (2017MNRAS.466.4826K; Kretschmer20), allowing detailed investigations on star formation activity within galaxies (Kraljic24).
@@
Furthermore, the corresponding increase in baryonic (and DM) mass resolution is also important, as poor resolutions introduce numerical biases (e.g., Klypin99; Ludlow19). Fine mass resolutions allow for the detection of structures with extremely low surface brightness, such as intracluster light (ICL), ultra-diffuse galaxies (UDGs), tidal streams, and ultra-faint dwarfs. A stellar particle mass of $\sim 10^{4}\,\,M_{\sun}$ enables the simulation to reach surface brightness levels as faint as $34\,{\rm mag\,arcsec^{-2}}$ at $z=0.1$. This is enough to reproduce results from deep imaging surveys such as Large Synoptic Survey Telescope (LSST), Euclid, and James Webb Space Telescope (JWST) (2022A&A...657A..92E).
@@
Therefore, more realistic and advanced cosmological simulations at the current stage require a sub-100 pc resolution and a high mass resolution. A major challenge arising from these requests is the high computational cost due to the enormous number of elements and associated calculations. A good bypass to satisfy the needs above and to avoid this difficulty is running a simulation for a small interested region with high resolutions, while the rest of the volume is treated with relatively poorer resolutions: that is, a zoom-in approach (e.g., Navarro94; Barnes17; CY17; Hopkins18b; Tremmel19; 2021A&A...651A.109D).
@@
The majority of recent high-resolution (sub-100 pc) zoom-in simulations focus on individual galaxies or average-density fields of the Universe. This has resulted in a lack of representative high-resolution counterparts for dense environments such as galaxy clusters. Galaxy clusters are thought to have formed from the growth of primordial over-density in the early Universe (2012ARA&A..50..353K), implying that developing clusters were once among the most active sites of star formation (2022ApJ...941....5J). Therefore, proto-clusters are the primary candidates for the sites for containing high-redshift galaxies that are frequently detected by the JWST (Labbe23; Carniani24).
@@
In this regard, we carry out the NewCluster simulation, a high-resolution cosmological hydrodynamic simulation designed to reproduce the formation and evolution of a galaxy cluster down to near-local redshift. NewCluster serves as a massive halo counterpart to existing high-resolution zoom-in simulations on average-density fields, NewHorizon and NewHorizon2, particularly because it shares many of its subgrid prescriptions with them. The simulation aims to monitor the formation and evolution of a massive cluster and galaxy evolution inside the extreme environment.
We expect morphological transformations of galaxies in clusters, which are responsible for the prevalence of early-type galaxies with old stellar populations (Dressler1980; Lewis02; Hogg04; Kauffmann04). This transition is thought to be triggered by the interplay of both internal and external processes. The energy and outflows released by stellar activity and central supermassive black holes are considered major factors in the quenching of star formation (2005ApJ...618..569M; 2011MNRAS.414..195T). The released gas forms the hot intracluster medium (ICM), which in turn drives an external quenching process for cluster members. Moving at very fast orbital velocity ($\sim 10^{3}\,{\rm km\,s^{-1}}$), galaxies experience rapid gas stripping through hydrodynamic ram pressure and DM and stellar stripping through gravitational tidal interactions (Smith16; Rhee17; Han18; Jung18). Such processes are likely also responsible for jellyfish (Chung09) and perhaps ultra-diffuse galaxies (vanDokkum2015). To closely trace the gravitational and hydrodynamic interactions that impact galaxy evolution, it is essential to achieve sufficient spatial resolution to resolve the kinematics and structural properties of galaxies. Such a level of detail was previously unavailable in simulations of galaxy clusters.
The impact of the cluster merger (starting at $z=0.8$) is also open to wild speculations. While NewCluster lacks the capability for statistical analysis, its high spatial resolution offers the opportunity to examine the merging process in greater detail, particularly regarding how member galaxies are affected by the global redistribution of the gravitational mass and ICM.
@@
In this paper, we summarise the results of the NewCluster simulation up to $z=0.8$, which corresponds to the half of the age of the universe and represents the latest available snapshot. At this redshift, a halo with a mass comparable to that of a galaxy cluster is present, which enables the simulation to reproduce the formation of an observed galaxy cluster analogue and its associated phenomena. This paper is organised as follows. In Sect. 2, we describe the simulation code, astrophysical prescriptions, and post-processing methods. In Sect. 3, we present a general overview of the simulation and the galaxy scaling relations at $z=0.8$, along with key examples of applications of the cluster simulation to various topics. In Sect. 4, we summarise the main findings of the paper.
@@ 2.1 Code: RAMSES-yOMP
RAMSES (2002A&A...385..337T) is an astrophysical simulation code based on the adaptive mesh refinement (AMR) technique. In this study, we use a specialised branch of the code, called RAMSES-yOMP (yomp), which has been extensively modified to enhance simulation performance. The improvement is primarily due to the implementation of a hybrid parallelization scheme that utilises a dual-level multiprocessing structure to enhance the parallel scalability, combining the message passing interface (MPI) and open multi-processing (OMP) libraries. A series of modifications to yOMP are described in yomp, including updates on subroutines such as the load-balancing scheme and the conjugate gradient Poisson solver, aimed at improving the overall performance of the simulation. Load balancing is enhanced by assigning greater computational weights to particles that require more processing, particularly those involved in complex feedback and density estimation. In a test suite simulating a cosmological volume with 1,536 computing cores, the modified code achieved approximately a two times speed-up compared to the pre-modified version.
@@ 2.2 Initial condition and zoom-in region
The initial condition is generated using MUSIC (2011MNRAS.415.2101H) at $z=50$ with the WMAP7 cosmology (2011ApJS..192...18K) with parameters $\Omega_{\Lambda}=0.728$, $\Omega_{m}=0.272$, $\sigma_{8}=0.810$, $n_{s}=0.967$, $h_{0}=0.703$ and $\Omega_{b}=0.0455$. The transfer function from 1998ApJ...496..605E and CAMB (2000ApJ...538..473L) is used to derive the initial distribution of DM and baryons, respectively. The zoom-in technique is utilised to efficiently simulate cluster regions in high resolution. First, we ran DM-only simulations of a box with a volume of $(100\,h^{-1}\,\mathrm{Mpc})^{3}$ with periodic boundary conditions. We use the Rockstar halo finder (2013ApJ...762..109B) to detect halos within the volume at $z=0$. Since the simulation is performed on only one galaxy cluster, we ran 15 DM-only simulations in total, and selected one target halo based on visual inspection. The target cluster whose evolution matches the following criteria was selected: 1. A cluster in a relaxed state at z = 0, without ongoing major mergers. 2. A cluster that had experienced a major merger during its assembly phase. A spherical region centred on the density peak of the target cluster is set with a radius of $3.5R_{\rm vir}$, where $R_{\rm vir}$ refers to the virial radius [^1] In Rockstar, the virial radius is defined the range within which contains $\Delta_{\rm vir}(z)\rho_{c}$ (Bryan1998; 2013ApJ...762..109B). of the target cluster, to record the identity of the confined particles and trace their initial distribution at $z=50$. The zoom-in region, corresponding to a $4.1\sigma$ overdensity above the Universal mean, is defined as a convex hull generated based on the distribution of the recorded particles in the initial volume. The selected convex hull roughly covers a comoving volume of $23,300\,\mathrm{Mpc}^{3}$ with a total mass of $8.6\times 10^{14}\,M_{\sun}$.
@@
The sudden increase in resolution at the specific scale factors may introduce numerical artifacts in the gas properties of galaxies. To mitigate this, we smooth out the transitions at which refinement is applied. We follow the approach of 2018MNRAS.477..983S, but additionally apply the same method to the Jeans length criterion. During the transition phase, we continuously varied the parameter $N_{\rm jeans}$ from 0 to 1 and $m_{\rm refine}$ from 100 to 8, at the transition phase to perform smooth refinement, i.e.,
@@
where $a$ is the current scale factor, $a_{0}$ is the centre of the transition phase (defined as 0.1, 0.2, 0.4, 0.8 at grid levels 18, 19, 20, 21), and $w$ is the half-width of the transition phase, set to $0.025$ in this work. $C$ is the refinement threshold and can be either $N_{\rm Jeans}$ or $m_{\rm refine}$, where $C_{L}$ and $C_{R}$ denote the initial and final values for the refinement transition, respectively. The parameter pair $C_{L},C_{R}]$ is set to $[0,1]$ for $N_{\rm Jeans}$, and $[100,8]$ for $m_{\rm refine}$. $m_{\rm refine}$ defines the mass increase required for a cell to be refined. For example, if $m_{\rm refine}=8$, the mass within a cell needs to be eight times higher than the initial cell mass to be eligible for refinement. $N_{\rm Jeans}$ is the parameter for the Jeans length refinement criterion, in which the cell with a hydrogen gas density $n\geq 5\,\rm H\,cm^{-3}$ can be refined if